OpenStudy (anonymous):

the sum of two numbers is 13 and their difference is 1. write a system of linear equations and solve the resulting system, let x be the larger number and y be the smaller number

OpenStudy (anonymous):

x+y = 13 x-y = 1 adding the equations together yields 2x=14 or x=7 x-y = 1 7-y=1 y=6

OpenStudy (anonymous):

thank you very much